Question 60 If M- $5,000, P= $20, and Q 2,000, then V is 2.0 4.0 6.0 8.0 Question 61
Answer #1

M * V = P * Q

5000 * V = 20 * 2000

V = 20 * 2000 / 5000

V = 8

Last option is correct